# Provable test-time adaptivity and distributional robustness of in-context learning

### v1

@@ F.2 Proof of Proposition 6
$=\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\biggr\}$
$=\mathbb{E}\biggl[\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}\biggr].$ (63)
Now conditional on $U^{(p)}$, we fix $g^{\prime}\in\mathcal{G}_{\alpha}^{(p)}$ such that $\bigl\|g^{\prime}\bigl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\bigr)-h^{\circ}\bigl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\bigr)\bigr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}$. We may assume without loss of generality that such $g^{\prime}$ exists, since otherwise the conditional probability in the expectation in (63) would be zero. Using the same argument as that in (51), we deduce that
$\mathbb{P}$ $\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
$\leq\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
@@
where $Q(U^{(p)})$ denotes the conditional distribution of $\frac{(U^{(p)})^{\top}X+\bm{1}_{p}}{2}$ given $U^{(p)}$, with $X\sim P_{X}$. By Lemma 27, for every $U\in V_{p}(\mathbb{R}^{d})$, there exists a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ such that $Q(U)(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$. We may thus apply (49), with $d$ replaced by $p$, $m$ replaced by $4m$ and $c_{0}$ replaced by $c_{0}^{\prime}$ therein, to deduce that
@@
Recall that $\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $\mathcal{G}_{\alpha,p}=\mathcal{G}_{\alpha}^{(p)}\circ\bigl\{x\mapsto\frac{U^{\top}x+\bm{1}_{p}}{2}:U\in V_{p}(\mathbb{R}^{d})\bigr\}$. Now let $\mathcal{N}_{1}$ be a $(\sigma\epsilon_{n,\alpha,p}/3)$-cover of $B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ with respect to $\|\cdot\|_{\infty}$ such that $|\mathcal{N}_{1}|\leq\exp(C_{1}n\epsilon_{n,\alpha,p}^{2})$ for some $C_{1}>0$, which exists by, e.g. Giné and Nickl (2021, Eq. (4.184)). Recall from (59) that for all $x,y\in[0,1]^{p}$,
@@
$\biggl|g\biggl($ $\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
$\leq\biggl|g\biggl(\frac{U^{\top}x+\bm{1}_{p}}{2}\biggr)-g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|+\biggl|g\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+\bm{1}_{p}}{2}\biggr)\biggr|$
@@ Appendix G Auxiliary lemmas
where the final inequality follows from the sub-Gamma assumption on $W(P,Z_{1})$. This shows that $\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$, so $n^{-1}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)$ is also sub-Gamma in the left tail with variance parameter $a\epsilon^{2}/n$ and scale parameter $c/n$ (e.g. Samworth and Shah, 2025+, Exercise 10.6.12). Hence, by (70),
@@
where the final inequality follows from Samworth and Shah (2025+, Proposition 10.6.8(a)). ∎
@@
The same bound holds for $W_{2}$, so by Samworth and Shah (2025+, Proposition 10.6.8(b)), $W_{2}$ is sub-Gamma in both tails with variance parameter $\frac{\epsilon^{2}}{2\sigma^{2}}$ and scale parameter $\frac{2R^{2}}{3\sigma^{2}}$. Finally, by Samworth and Shah (2025+, Exercise 10.6.12), we have that $W=W_{1}+W_{2}$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c=\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
@@
Let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ that satisfies the assumptions of Lemma 21. Let $p\in[d]$, $U\in V_{p}(\mathbb{R}^{d})$, $X\sim P_{X}$ and let $Q$ denote the distribution of $\frac{U^{\top}X+\bm{1}_{p}}{2}$. Then there exist $c_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p)$ and a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ with the property that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$.
@@
$Q\biggl(\frac{U^{\top}A_{0}+\bm{1}_{d}}{2}\biggr)=\mathbb{P}(X\in A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0}).$
Moreover, $(U^{\top}A+\bm{1}_{d})/2$ is a Euclidean ball of radius $\tau/2$, so it contains a hypercube of side length $\frac{\tau}{2\sqrt{p}}$. Now suppose that $p\in[d-1]$. For $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$ and $1\leq j_{1}\leq j_{2}\leq d$, we write $v_{j_{1}:j_{2}}\coloneqq(v_{j_{1}},\ldots,v_{j_{2}})^{\top}\in\mathbb{R}^{j_{1}-j_{2}+1}$. Let $o\in\mathbb{R}^{d}$ be the centre of $A$ and let $K\subseteq A$ be the Euclidean ball of radius $\tau/2$ and centred at $o$. We may write $U=VP$ where $V\in\mathbb{R}^{d\times d}$ is orthogonal and $P\in\mathbb{R}^{d\times p}$ is the first $p$ columns of $I_{d}$. Let $S\subseteq K$ be Borel measurable, and define $T\coloneqq\{x\in V^{\top}A:x_{1:p}\in P^{\top}V^{\top}S\}$. Then $VT\subseteq A$; moreover, if $y=(V^{\top}x)_{1:p}$ for some $x\in S\subseteq K$ and $z\in\mathbb{R}^{d-p}$ satisfies $\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2$, then by the choice of $K$,
@@
$Q\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr)$ $=\mathbb{P}\biggl(\frac{P^{\top}V^{\top}X+\bm{1}_{p}}{2}\in\frac{P^{\top}V^{\top}S+\bm{1}_{p}}{2}\biggr)$
@@
$=c_{0}v(\tau,d,p)\mathrm{Vol}_{p}(P^{\top}V^{\top}S)\eqqcolon c_{0}^{\prime}\mathrm{Vol}_{p}\biggl(\frac{U^{\top}S+\bm{1}_{p}}{2}\biggr).$
Since this holds for all measurable $S\subseteq K$, we deduce that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all $A_{0}^{\prime}\subseteq\frac{U^{\top}K+\bm{1}_{p}}{2}$. Now $\frac{U^{\top}K+\bm{1}_{p}}{2}$ is a ball of radius $\tau/4$ in $\mathbb{R}^{p}$, so it contains a hypercube $A^{\prime}$ of side length $\frac{\tau}{2\sqrt{p}}$. ∎

### v2

@@ F.2 Proof of Proposition 6
$=\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\biggr\}$
$=\mathbb{E}\biggl[\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}\biggr].$ (65)
Now conditional on $U^{(p)}$, we fix $g^{\prime}\in\mathcal{G}_{\alpha}^{(p)}$ such that $\bigl\|g^{\prime}\bigl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\bigr)-h^{\circ}\bigl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\bigr)\bigr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}$. We may assume without loss of generality that such $g^{\prime}$ exists, since otherwise the conditional probability in the expectation in (65) would be zero. Using the same argument as that in (53), we deduce that
$\mathbb{P}$ $\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1}{2}\biggr)-h^{\circ}\biggl(\frac{(U^{\circ})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
$\leq\mathbb{P}\biggl\{\biggl\|\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr)\biggr\|_{L^{2}(P_{X})}^{2}\leq 4c_{0}^{\prime}m^{2}n^{-2\alpha/(2\alpha+p)}\,\bigg|\,U^{(p)}\biggr\}$
@@
where $Q(U^{(p)})$ denotes the conditional distribution of $\frac{(U^{(p)})^{\top}X+1_{p}}{2}$ given $U^{(p)}$, with $X\sim P_{X}$. By Lemma 27, for every $U\in V_{p}(\mathbb{R}^{d})$, there exists a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ such that $Q(U)(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$. We may thus apply (51), with $d$ replaced by $p$, $m$ replaced by $4m$ and $c_{0}$ replaced by $c_{0}^{\prime}$ therein, to deduce that
@@
Recall that $\mathcal{G}_{\alpha}^{(p)}\subseteq B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ and $\mathcal{G}_{\alpha,p}=\mathcal{G}_{\alpha}^{(p)}\circ\bigl\{x\mapsto\frac{U^{\top}x+1_{p}}{2}:U\in V_{p}(\mathbb{R}^{d})\bigr\}$. Now let $\mathcal{N}_{1}$ be a $(\sigma\epsilon_{n,\alpha,p}/3)$-cover of $B_{\infty,\infty}^{\alpha}([0,1]^{p},2C_{0})$ with respect to $\|\cdot\|_{\infty}$ such that $|\mathcal{N}_{1}|\leq\exp(C_{1}n\epsilon_{n,\alpha,p}^{2})$ for some $C_{1}>0$, which exists by, e.g. Giné and Nickl (2021, Eq. (4.184)). Recall from (61) that for all $x,y\in[0,1]^{p}$,
@@
$\biggl|g\biggl($ $\frac{U^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|$
$\leq\biggl|g\biggl(\frac{U^{\top}x+1_{p}}{2}\biggr)-g\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|+\biggl|g\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)-g^{\prime}\biggl(\frac{(U^{\prime})^{\top}x+1_{p}}{2}\biggr)\biggr|$
@@ Appendix G Auxiliary Lemmas
where the final inequality follows from the sub-Gamma assumption on $W(P,Z_{1})$. This shows that $\int_{B}W(P,Z_{1})\,\mathrm{d}\Pi(P)$ is sub-Gamma in the left tail with variance parameter $a\epsilon^{2}$ and scale parameter $c$, so $n^{-1}\sum_{i=1}^{n}\int_{B}W(P,Z_{i})\,\mathrm{d}\Pi(P)$ is also sub-Gamma in the left tail with variance parameter $a\epsilon^{2}/n$ and scale parameter $c/n$. Hence, by (72),
@@
where the final inequality follows from Bernstein’s inequality (Boucheron, Lugosi and Massart, 2013, Corollary 2.11). ∎
@@
The same bound holds for $W_{2}$, so by Boucheron, Lugosi and Massart (2013, Theorem 2.10), $W_{2}$ is sub-Gamma in both tails with variance parameter $\frac{\epsilon^{2}}{2\sigma^{2}}$ and scale parameter $\frac{2R^{2}}{3\sigma^{2}}$. Finally, it follows that that $W=W_{1}+W_{2}$ is sub-Gamma in both tails with variance parameter $\bigl(8\vee\frac{2}{\sigma^{2}}\bigr)\epsilon^{2}$ and scale parameter $c=\frac{2R}{\sigma}\vee\frac{4R^{2}}{3\sigma^{2}}$.
@@
Let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ that satisfies the assumptions of Lemma 21. Let $p\in[d]$, $U\in V_{p}(\mathbb{R}^{d})$, $X\sim P_{X}$ and let $Q$ denote the distribution of $\frac{U^{\top}X+1_{p}}{2}$. Then there exist $c_{0}^{\prime}\equiv c_{0}^{\prime}(c_{0},\tau,d,p)$ and a hypercube $A^{\prime}\subseteq[0,1]^{p}$ of side length $\frac{\tau}{2\sqrt{p}}$ with the property that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all measurable $A_{0}^{\prime}\subseteq A^{\prime}$.
@@
$Q\biggl(\frac{U^{\top}A_{0}+1_{d}}{2}\biggr)=\mathbb{P}(X\in A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0}).$
Moreover, $(U^{\top}A+1_{d})/2$ is a Euclidean ball of radius $\tau/2$, so it contains a hypercube of side length $\frac{\tau}{2\sqrt{p}}$. Now suppose that $p\in[d-1]$. For $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$ and $1\leq j_{1}\leq j_{2}\leq d$, we write $v_{j_{1}:j_{2}}\coloneqq(v_{j_{1}},\ldots,v_{j_{2}})^{\top}\in\mathbb{R}^{j_{1}-j_{2}+1}$. Let $o\in\mathbb{R}^{d}$ be the centre of $A$ and let $K\subseteq A$ be the Euclidean ball of radius $\tau/2$ and centred at $o$. We may write $U=VP$ where $V\in\mathbb{R}^{d\times d}$ is orthogonal and $P\in\mathbb{R}^{d\times p}$ is the first $p$ columns of $I_{d}$. Let $S\subseteq K$ be Borel measurable, and define $T\coloneqq\{x\in V^{\top}A:x_{1:p}\in P^{\top}V^{\top}S\}$. Then $VT\subseteq A$; moreover, if $y=(V^{\top}x)_{1:p}$ for some $x\in S\subseteq K$ and $z\in\mathbb{R}^{d-p}$ satisfies $\|z-(V^{\top}o)_{(p+1):d}\|_{2}\leq\tau/2$, then by the choice of $K$,
@@
$Q\biggl(\frac{U^{\top}S+1_{p}}{2}\biggr)$ $=\mathbb{P}\biggl(\frac{P^{\top}V^{\top}X+1_{p}}{2}\in\frac{P^{\top}V^{\top}S+1_{p}}{2}\biggr)$
@@
$=c_{0}v(\tau,d,p)\mathrm{Vol}_{p}(P^{\top}V^{\top}S)\eqqcolon c_{0}^{\prime}\mathrm{Vol}_{p}\biggl(\frac{U^{\top}S+1_{p}}{2}\biggr).$
Since this holds for all measurable $S\subseteq K$, we deduce that $Q(A_{0}^{\prime})\geq c_{0}^{\prime}\mathrm{Vol}_{p}(A_{0}^{\prime})$ for all $A_{0}^{\prime}\subseteq\frac{U^{\top}K+1_{p}}{2}$. Now $\frac{U^{\top}K+1_{p}}{2}$ is a ball of radius $\tau/4$ in $\mathbb{R}^{p}$, so it contains a hypercube $A^{\prime}$ of side length $\frac{\tau}{2\sqrt{p}}$. ∎
Appendix H ICL Example
When we provide ChatGPT with the first two in-context examples and the query from the right column of Figure 1 (i.e. the prompt is ‘Cat $\to$ 10, Fish $\to$ 25, Bird $\to$ ?’), we obtain the following output (though this may vary due to uncertainty in the sampling procedure):
There isn’t a unique answer from just Cat $\to$ 10 and Fish $\to$ 25. With only two example pairs, infinitely many rules can fit them, and those rules would give different values for Bird. So Bird $\to$ ? is underdetermined (cannot be determined uniquely). If you give one more example mapping (any other word $\to$ number), I can pin down the rule and compute Bird.
When ChatGPT is provided with all four examples and the query from the right column of Figure 1 (i.e. now the prompt is ‘Cat $\to$ 10, Fish $\to$ 25, Deer $\to$ 17, Dog $\to$ 13, Bird $\to$ ?’), we obtain the following:
The pattern is: Number = (number of letters in the word) $\times$ (alphabet position of the first letter) + 1 Check: Cat $\to$ 3 $\times$ C(3) + 1 = 10 Fish $\to$ 4 $\times$ F(6) + 1 = 25 Deer $\to$ 4 $\times$ D(4) + 1 = 17 Dog $\to$ 3 $\times$ D(4) + 1 = 13 So: Bird $\to$ 4 $\times$ B(2) + 1 = 9
Appendix I Further Details for Simulations in Section 5
For training, we used AdamW (Loshchilov and Hutter, 2019), with weight decay 0.001 and learning rate 0.0003 (with a cosine annealing schedule). We also employed pre-layer normalisation (Xiong et al., 2020) and gradient clipping to stabilise the optimisation. Our Transformer was trained for 200,000 optimisation steps. At each step, we sampled 50 independent single-index regression functions, and for each function, we randomly generated 12 prompts and queries (i.e. the covariate vectors and queries are random while keeping the regression function fixed), yielding a total of 600 pretraining sequences per optimisation step. The training was done on 12 NVIDIA A100 GPUs (80GB) over approximately 17 hours. To facilitate training, the difficulty of the tasks was progressively increased via a curriculum learning strategy. In the initial warm-up phase (the first 10,000 training steps), the ambient dimension $d$ was gradually increased from 1 to 5, while keeping $p=1$ and $\alpha=4$ fixed. Subsequently, for the next 50,000 steps, we introduced a smoothness curriculum, where regression functions were drawn from increasingly diverse mixtures whose smoothness parameters were gradually decreased over time. Finally, for the remaining 140,000 steps, training proceeded with a uniform mixture over tasks, where the smoothness parameter was sampled uniformly from $\{2,\,2.5,\,3,\,4\}$.